1. Find the domain of the function 𝑓(π‘₯)=ln(βˆ’2π‘₯2βˆ’π‘₯βˆ’6)+√π‘₯2βˆ’1.

Β 

2. Find the inverse function of the function 𝑓(π‘₯)=π‘₯2βˆ’4π‘₯+5,π‘₯βˆˆγ€ˆ3,4⟩. Find the domain and the range of the inverse function.

Β 

3. Construct the tangent line to the graph of the function 𝑓(π‘₯)=4π‘₯β‹…βˆšπ‘₯βˆ’2β‹…βˆšπ‘₯ which is parallel to the line 𝑦=π‘₯.

Β 

4. Find the maximal intervals of monotonicity of the function 𝑓(π‘₯)=𝑒π‘₯+3π‘₯2+2π‘₯+6.

Β 

5. Find the integral ∫6β‹…π‘₯3⋅𝑒π‘₯2+2𝑑π‘₯.

Β 

6. Find the general solution of the differential equation π‘₯2+1+𝑦′⋅cos(𝑦)=0.Β 


1
Expert's answer
2022-05-26T07:03:25-0400

1.

a)


"f(x)=\ln(-2x^2-x-6)+\sqrt{x^2-1}"


"-2x^2-x-6>0""x^2-1\ge0"


"D=(-1)^2-4(-2)(-6)=-47<0"

Then "-2x^2-x-6<0, x\in \R"


"Domain:\{\}"


b)


"f(x)=\ln(-\dfrac{2}{x^2-x-6})+\sqrt{x^2-1}"


"x^2-x-6<0""x^2-1\ge0"

"(x+2)(x-3)<0"

"x\le -1\ or\ x\ge1"

Domain:"(-2, -1]\cup[1,3)"


2.


"𝑓(𝑥)=x^2−4x+5,x\in [3, 4]"

"x_v=-\dfrac{-4}{2(1)}=2"

The function "f" increases on "(3, 4)"


"f(3)=(3)^2\u22124(3)+5=2"

"f(4)=(4)^2\u22124(4)+5=5"

Domain: "[3, 4]"

Range: "[2, 5]"


"y=x^2-4x+5, 3\le x\le4"

ChangeΒ "x"Β andΒ "y"

"x=y^2-4y+5, 3\le y\le 4"

Solve forΒ "y"

"y^2-4y+4=x-1""(y-2)^2=x-1"

Since "3\le y\le 4"


"y-2=\sqrt{x-1}"




"y=2+\sqrt{x-1}"

Then


"f^{-1}(x)=2+\sqrt{x-1}"


Domain: "[2, 5]"

Range: "[3, 4]"


3.


"f(x)=4x\sqrt{x}-2\sqrt{x}"

Domain: "[0, \infin)"


"f'(x)=4(\dfrac{3}{2})\sqrt{x}-\dfrac{2}{2\sqrt{x}}=\dfrac{6x-1}{\sqrt{x}}"

"slope=f'(x)=\dfrac{6x-1}{\sqrt{x}}=1"

"6x-\sqrt{x}-1=0"

"(3\sqrt{x}+1)(2\sqrt{x}-1)=0"

Since "\sqrt{x}\ge0," we take "2\sqrt{x}-1=0"


"\sqrt{x}=\dfrac{1}{2}"

"x=\dfrac{1}{4}"

"f(\dfrac{1}{4})=4(\dfrac{1}{4})\sqrt{\dfrac{1}{4}}-2\sqrt{\dfrac{1}{4}}=-\dfrac{1}{2}"

The tangent line to the graph is


"y+\dfrac{1}{2}=x-\dfrac{1}{4}"


"y=x-\dfrac{3}{4}"

4.


"f(x)=e^x+3x^2+2x+6"

Domain: "(-\infin, \infin)"


"f'(x)=e^x+6x+2"

"f'(x)=0=>e^x+6x+2=0"

"(e^x+6x+2)'=e^x+6>0, x\in \R"

The function "g(x)=e^x+6x+2" is strictly increasing on "(-\infin, \infin)."

The equation "e^x+6x+2=0" has the only solution "x\approx-0.4406."


The function "f(x)" is monotonic decreasing on "(-\infin, -0.4406)."

The function "f(x)" is monotonic increasing on "( -0.4406, \infin)."


5.


"\int 6x^3e^{x^2+2}dx"

"t=x^2, dt=2xdx"


"\int 6x^3e^{x^2+2}dx=3e^2\int te^tdt"

"\int te^t dt"

"u=t, du=dt"

"dv=e^tdt, v=e^t"

"\int te^t dt=te^t -\int e^t dt=te^t-e^t+C_1"


"\int 6x^3e^{x^2+2}dx=3e^{x^2+2}(x^2-1)+C"

Β 6.


"x^2+1+y'\cos(y)=0"

"\cos(y)dy=-(x^2+1)dx"

Integrate


"\int \cos(y)dy=-\int (x^2+1)dx"

"\sin (y)=-\dfrac{x^3}{3}-x+C"

"\sin (y)+\dfrac{x^3}{3}+x=C"